### v4

@@ 2.2.1 Layer construction with ancilla qubits
There are 4 hole pockets, each of fractional area $p/8$, which are obtained from the hybridization of the $c_{\alpha}$ band of density $1+p$ and the $f_{1\alpha}$ band of density 1 in Eq. (2.4). (In contrast, in spin density wave theory, we would hybridize the $c_{\alpha}$ band of density $1+p$, with a $(\pi,\pi)$ momentum-shifted $c_{\alpha}$ band also of density $1+p$ to obtain pockets of area $p/4$.) The pockets of Fig. 11 do have faint, but visible, ‘backsides’. In contrast, the observations show only ‘Fermi arcs’ corresponding to the front sides of the pockets (See Fig. 24a later), with no visible backsides. We will consider the influence of fluctuations on this photoemission spectrum in Section 5.1, where we will see that they do indeed produce the observed Fermi arcs.
@@
More recently, compelling evidence for hole pockets has emerged at low doping in the higher temperature pseudogap metal phase [17, 18]. Chan et al [18] have observed the remarkable Yamaji effect, which allows deduction of pocket area only from a knowledge of the layer spacing and the value of the Yamaji angle at which there is peak in the magnetoresistance. At these higher temperatures there is no ordering, and the Hall effect is positive, indicating the absence of electron pockets. Their results are consistent with the FL* prediction of hole pockets of area $p/8$ [36, 37, 52, 151], as we noted earlier near Fig. 6. This is to be constrasted with pockets of area $p/4$ in the fluctuating spin density wave state [32, 33, 34, 35]. The Yamaji angle observation of Chan et al. [18] is therefore the long-sought direct signature of fractionalization in the cuprates.
@@ 3 Critical spin liquid on the square lattice
Our FL* theory of the pseudogap has led us to a mean-field theory described by the mean-field Kondo lattice Hamiltonian in Eq. (2.4), and a decoupled spin liquid of ${\bm{S}}_{2}$ spins. The next step requires us to couple these sectors to each other in the ALM. In the present section, we consider the physics of the ${\bm{S}}_{2}$ layer alone by discussing the theory of spin liquids on the square lattice. We chose a spinon Fermi surface for the spin liquid of the ${\bm{S}}_{1}$ spins in Section 2.2 to obtain the hole pockets by hybridizing it to the large Fermi surface of the top layer. The power of the ALM method is that it gives us the flexibility to make a different choice for the spin liquid of the ${\bm{S}}_{2}$ spins—we will choose it from studies of insulating square lattice antiferromagnets.
@@ 3.2 Fermionic partons
Next we insert Eq. (3.32) into Eq. (3.1), and perform Hubbard-Stratonovich transformation to obtain an effective Hamiltonian for the spinons, following the same procedure as for bosonic spinons. We skip the intermediate steps, and focus directly on the fermion bilinear Hamiltonian on symmetry grounds. From the gauge transformations in Eq. (3.33), and the global spin rotation in Eq. (3.34), we anticipate a spinon hopping term of the form
$\mbox{Tr}\left(\mathcal{F}_{{\bm{i}}}^{\dagger}U_{{\bm{i}}{\bm{j}}}^{\phantom{\dagger}}\mathcal{F}_{{\bm{j}}}^{\phantom{\dagger}}\right)$ (3.36)
which is invariant under both transformations. Here we have introduced a SU(2) gauge field $U_{{\bm{i}}{\bm{j}}}=U_{{\bm{j}}{\bm{i}}}^{\dagger}$ on each lattice link upon which the SU(2) gauge transformation acts as
However, the identity
$\mbox{Tr}\left(\mathcal{F}_{{\bm{i}}}^{\dagger}\mathcal{F}_{{\bm{j}}}^{\phantom{\dagger}}\right)=-\mbox{Tr}\left(\mathcal{F}_{{\bm{j}}}^{\dagger}\mathcal{F}_{{\bm{i}}}^{\phantom{\dagger}}\right)$ (3.38)
implies that we need a pure-imaginary hopping in a hermitian Hamiltonian in the mean-field with $U_{{\bm{i}}{\bm{j}}}=1$. So we have the mean-field nearest-neighbor spin liquid Hamiltonian for the spinons of the $\pi$-flux phase [64]:
@@
If we had not used the pure imaginary hopping in Eq. (3.41), then the mean-field Hamiltonian would break (‘Higgs’) the SU(2) gauge symmetry to a smaller symmetry. A ‘staggered flux’ ansatz which breaks the SU(2) down to U(1) has commonly been used in the literature [57]. However, it is now known that this U(1) spin liquid allows single monopole perturbations [157, 158] (unlike the quadrupole monopole perturbations in Eq. (3.27)), and such single monopole terms are expected to drive a strong instability to confinement. So we don’t consider this U(1) spin liquid here.
@@
Going beyond mean field theory, while still remaining on the lattice, we extend the mean field Hamiltonian in Eq. (3.41) by adding the gauge field $U_{ij}$ as in Eq. (3.36). Then, by gauge invariance, we extend Eq. (3.41) to
@@ 3.4 Direct observation of spinons
For the doped system, it has recently been argued [89, 114] that RIXS measurements [164, 165] do indeed provide support for the presence of spinons. At half-filling, there is sharp spin wave excitation extending to high energy $\sim 400$meV. This excitation survives upon doping even at high energies, but becomes much broader. This broad spectrum has similarities to that expected from the $\pi$-flux spin liquid, as computed near the destruction of the Néel order in the insulating $J_{1}$-$J_{2}$ model by Ref. [166]. See Figs. 8 and 9 in Ref. [114] for a comparison.
@@ 4 Confinement to $d$-wave superconductivity and charge order at half-filling
In terms of adiabatic continuity, this $d$-wave superconductor is precisely the superconductor observed in the cuprates. However, the $d$-wave superconductor obtained in this section has one significant quantitative difference from the observations: it has a Lorentz-invariant form of its dispersion, with the two velocities only the square lattice diagonals, $v_{F}$ and $v_{\Delta}$, being equal to each other (see Fig. 26B). The cuprates instead have $v_{F}\gg v_{\Delta}$. We will resolve this problem in an interesting manner in Section 6.1 when we consider the transition from FL* to a $d$-wave superconductor in the doped case.
@@
A boson with the same quantum numbers as our $B$ has been considered in earlier work [66, 47, 81], but with an important difference. In the earlier work, the expectation value of $B^{\dagger}B$ was set to be equal to the doping density. That is not the case in our work, as the doping density also includes the density of fermionic holes (see Section 2). In this section, we are considering the particle-hole symmetric case at half-filling, and $\langle B^{\dagger}B\rangle$ is non-zero even though there is no doping (this is similar to the situation in the superfluid-insulator transition of bosons [180]).
@@
These considerations lead to the energy functional $\mathcal{E}_{2}[B,U]+\mathcal{E}_{4}[B,U]$ with terms quadratic and quartic in $B$ respenctively:
@@ 5 SU(2) gauge theory of the cuprate pseudogap
specified in Eqs. (2.4), (3.47), (5.3), (4.17). This Hamiltonian has 3 fermions $c_{\alpha}$, $f_{1\alpha}$, $f_{\alpha}$ whose transformations under SU(2) gauge, spin rotation, and electromagnetic charge symmetries are in Eqs. (4.8), (4.9), (4.10), with $f_{1\alpha}$ transforming just like $c_{\alpha}$. As we noted earlier, the boson $\Phi$ will be treated as a $c$-number constant for now, although we will consider its quantum and thermal fluctuations in Section 8 when we consider the transition to the FL phase at large doping. The completion of $\mathcal{H}_{\rm pseudogap}$ for the case of dynamic $\Phi$ is presented in the Appendix.
@@ 5.1 Photoemission spectrum
In the mean-field analysis of Section 2.2.4, we obtained hole pocket description of the Fermi surface in Fig. 11, along with the dispersion antinodal pseudogap in Fig. 12. Here, we will consider modification of the Fermi surface from the coupling between the bottom two layers of the ALM. The computations of Pandey et al [50] included the coupling to the spinons in $\mathcal{H}_{\rm pseudogap}$ in Eq. (5.4). This can be expected to have a stronger effect along the Brillouin zone diagonals proximate to the spinon nodes, as shown in Fig. 22. Indeed they found that the backside of the hole pockets are suppressed, and spectral weight turned into Fermi arcs, similar to observations.
@@
Figure 23: From Ref. [50]. Mean-Field phase diagram of Eq. (4.17) a function of the further neighbor interactions $V_{11}$ and $V_{22}$, extending that in Fig. 18. The other parameters are r=0.732r=-0.732, $w=0.40$, $u=0$, $V_{1}=0$, $g=0.021446$, $J_{1}=K_{1}={2}/({4(1+\sqrt{2})^{2}})$. ddSC is $d$-wave superconductivity, CDWn×m is a charge density wave with a supercell with $n\times m$ lattice sites. The red cross marks the parameter values chosen for the Monte Carlo simulations of Eq. (5.5).
@@
Figure 24: The right panel shows the electron spectral function measured in photoemission computed in Ref. [50] from $\mathcal{H}_{\rm KLmf}+\mathcal{H}_{SLf}+\mathcal{H}_{Y}$ after averaging over the Monte Carlo simulation of $B$ and $U_{{\bm{i}}{\bm{j}}}$ in Eq. (5.5). The spectrum with $B=U=0$ is in Fig. 22, showing both the hole pockets and the spinons. The left panel is experimental photoemission data from Ref. [188].
@@ 5.2 Magnetotransport
Given the role of thermal $B$ fluctuations in removing the pocket backsides in the photoemission spectrum in Fig. 24, it is legitimate to ask if the thermal $B$ fluctuations will also wash out quantum oscillations. The latter probe does not eject electrons form the sample, unlike photoemission, and so its behavior can be distinct. This question was investigated by Pandey et al. [50]: they averaged the quantum oscillations associated with $\mathcal{H}_{\rm KLmf}+\mathcal{H}_{SLf}+\mathcal{H}_{Y}$ over gaussian fluctations of $B$. They found that the quantum oscillations are damped by thermal fluctuations but survive in regimes where the Fermi arc spectrum in Fig. 24 was well established, as shown in Fig. 25.
@@ 6 From the pseudogap to $d$-wave superconductivity and competing orders
This section addresses the fate of the FL* pseudogap as the temperature is lowered, upon including the coupling to the ${\bm{S}}_{2}$ spin liquid in $\mathcal{H}_{\rm pseudogap}$ in Eq. (5.4). This Hamiltonian specifically chose the $\pi$-flux spin liquid of Section 3.2 for the ${\bm{S}}_{2}$ layer. We will show that for this spin liquid there is a transition to a conventional BCS-type $d$-wave superconductor, with anisotropic nodal velocities for the Boboliubov quasiparticles, and $h/(2e)$ vortices. Nevertheless the transition itself is not of the BCS type with a Cooper-pairing instability of a Fermi surface. Instead, the transition is driven by the confinement of the fractionalized excitations of the ${\bm{S}}_{2}$ spin liquid. We also find nearby instabilities to charge ordering, consistent with observations.
@@ 6.1 Anisotropic velocities in the $d$-wave superconductor
We now show how the problem of isotropic quasiparticle velocities, noted in Section 4, is resolved by the presence of the pocket Fermi surfaces described by $\mathcal{H}_{\rm KLmf}$ in Eq. (2.4). The discussion below is based on the detailed computations presented in Refs. [84, 85].
@@ 7 The SYK model
We now turn to a different critical quantum liquid, associated with mobile electrons, the SYK model. This is a zero-dimensional model for which the absence of quasiparticle excitations is well established. A direct and extensive application of the zero-dimensional SYK model has been to the low energy quantum theory of charged black holes, and this is reviewed elsewhere [195]. There are also connections to experiments on graphene flakes [196]. But our primary interest here will be extensions to two-dimensional models relevant to the strange metal state of the cuprates and other correlated electron materials—we will turn to this in Section 8.
@@ 8 From the SYK model to strange metals
    where $n$ is the known conduction electron density, and $m^{\ast}$ is an electronic effective mass. In some experiments, the transport mass $m_{\rm trans}^{\ast}$ of Eq. (8.2) is used in Eq. (8.4), while other experiments use the $m^{\ast}$ determined from thermodynamic measurements. In the form Eq. (8.4), the absolute value of $\tau_{\rm trans}$ can be deduced from experimental observations. In the strange metal, such a value is found to obey ‘Planckian’ behavior with [229, 230, 231]
    $\frac{1}{\tau_{\rm trans}}=\alpha\,\frac{k_{B}T}{\hbar}\,$, (8.5)
with $\alpha$ a numerical constant of order unity. Measurements of $1/\tau_{\rm trans}$ in La1.6-xNd0.4SrxCuO4 in angle-dependent magnetotransport show $\alpha=1.2\pm 0.4$ [231] upon using the thermodynamic $m^{\ast}$.

### v6

@@ 2.2.1 Layer construction with ancilla qubits
There are 4 hole pockets, each of fractional area $p/8$, which are obtained from the hybridization of the $c_{\alpha}$ band of density $1+p$ and the $f_{1\alpha}$ band of density 1 in Eq. (2.4). (In contrast, in spin density wave theory, we would hybridize the $c_{\alpha}$ band of density $1+p$, with a $(\pi,\pi)$ momentum-shifted $c_{\alpha}$ band also of density $1+p$ to obtain pockets of area $p/4$.) The pockets of Fig. 11 do have faint, but visible, ‘backsides’. In contrast, the observations show only ‘Fermi arcs’ corresponding to the front sides of the pockets (see Fig. 24a later), with no clearly visible backsides (although indications of backsides have been discussed [193, 30, 194]). We will consider the influence of thermal fluctuations on this photoemission spectrum in Section 5.1, where we will see that they can indeed produce the observed Fermi arcs in the intermediate temperature pseudogap.
@@
More recently, compelling evidence for hole pockets has emerged at low doping in the higher temperature pseudogap metal phase [9, 10]. Chan et al [10] have observed the remarkable Yamaji effect.
The Yamaji effect requires coherent interlayer transport of the charge carriers, and this already places stringent constraints on the model of the pseudogap, and favoring FL*. The holon metal does not have coherent interlayer transport, while the fluctuating spin density wave state [50, 51, 52, 53] does not have significant interlayer spin correlations [80, 81]. needed for coherent transport between SDW pockets.
Moreover, the Yamaji effect allows deduction of pocket area only from a knowledge of the layer spacing and the value of the Yamaji angle at which there is peak in the magnetoresistance. At these higher temperatures there is no ordering, and the Hall effect is positive, indicating the absence of electron pockets. Their results are consistent with the FL* prediction of hole pockets of area $p/8$ [12, 13, 69, 2, 189, 78, 79] which can tunnel coherently between layers, as we noted earlier near Fig. 6. This is to be contrasted with pockets of area $p/4$ in the SDW state.
The Yamaji angle observation of Chan et al. [10] is therefore the long-sought direct signature of fractionalization in the cuprates.
@@ 3 Critical spin liquid on the square lattice
Our FL* theory of the pseudogap has led us to a mean-field theory described by the mean-field Kondo lattice Hamiltonian in Eq. (2.4), and a decoupled spin liquid of ${\bm{S}}_{2}$ spins in the bottom layer. The next step requires us to couple these sectors to each other in the ALM. In the present section, we consider the physics of the ${\bm{S}}_{2}$ layer alone by discussing the theory of spin liquids on the square lattice. We chose a spinon Fermi surface for the spin liquid of the ${\bm{S}}_{1}$ spins in the middle layer in Section 2.2 to obtain the hole pockets by hybridizing it to the large Fermi surface of the top layer. The power of the ALM method is that it gives us the flexibility to make a different choice for the spin liquid of the ${\bm{S}}_{2}$ spins in the bottom layer—we will choose it from studies of insulating square lattice antiferromagnets.
@@ 3.2 Fermionic partons
Next we insert Eq. (3.32) into Eq. (3.1), and perform Hubbard-Stratonovich transformation to obtain an effective Hamiltonian for the spinons, following the same procedure as for bosonic spinons. We skip the intermediate steps, and focus directly on the fermion bilinear Hamiltonian on symmetry grounds; we obtain the mean-field nearest-neighbor spin liquid Hamiltonian for the spinons of the $\pi$-flux phase [94]:
@@
In the chosen form, it is evident that $\mathcal{H}_{Slf}$ is invariant under the global spin rotation in Eq. (3.34), and also at least the gobal gauge transformation associate with Eq. (3.33). We have chosen the hopping to be pure imaginary because of the identity
$\mbox{Tr}\left(\mathcal{F}_{{\bm{i}}}^{\dagger}\mathcal{F}_{{\bm{j}}}^{\phantom{\dagger}}\right)=-\mbox{Tr}\left(\mathcal{F}_{{\bm{j}}}^{\dagger}\mathcal{F}_{{\bm{i}}}^{\phantom{\dagger}}\right)\,.$ (3.40)
@@
Going beyond mean field theory, while still remaining on the lattice, we extend the mean field Hamiltonian in Eq. (3.38) to be invariant under space-dependent SU(2) gauge symmetries. To achieve this, we need to introduce a SU(2) gauge field $U_{{\bm{i}}{\bm{j}}}=U_{{\bm{j}}{\bm{i}}}^{\dagger}$ on each lattice link upon which the SU(2) gauge transformation acts as
Then, by gauge invariance, we extend Eq. (3.38) to
@@
If we had not used the pure imaginary hopping in Eq. (3.38), then the mean-field Hamiltonian would break (‘Higgs’) the SU(2) gauge symmetry to a smaller symmetry. A ‘staggered flux’ ansatz which breaks the SU(2) down to U(1) has commonly been used in the literature [8]. However, it is now known that this U(1) spin liquid allows single monopole perturbations [198, 199] (unlike the quadrupole monopole perturbations in Eq. (3.27)), and such single monopole terms are expected to drive a strong instability to confinement. So we don’t consider the staggered-flux U(1) spin liquid.
@@ 3.4 Direct observation of spinons
For the doped system, it has recently been argued [118, 121] that RIXS measurements [205, 206] do indeed provide support for the presence of spinons. At half-filling, there is sharp spin wave excitation extending to high energy $\sim 400$meV. This excitation survives upon doping even at high energies, but becomes much broader. This broad spectrum has similarities to that expected from the $\pi$-flux spin liquid, as computed near the destruction of the Néel order in the insulating $J_{1}$-$J_{2}$ model by Ref. [207]. See Figs. 8 and 9 in Ref. [121] for a comparison. Ref. [208] has considered the influence of charge order on the fermionic spinons, and shown that it can reproduce the ‘hourglass’ structure of the incommensurate spin corrections.
@@ 4 Confinement to $d$-wave superconductivity and charge order at half-filling
In terms of adiabatic continuity, this $d$-wave superconductor is precisely the superconductor observed in the cuprates. However, the $d$-wave superconductor obtained in this section has one significant quantitative difference from the observations, noted in Section 1: it has a Lorentz-invariant form of its dispersion, with the two velocities only the square lattice diagonals, $v_{F}$ and $v_{\Delta}$, being equal to each other (see Fig. 26B). The cuprates instead have $v_{F}\gg v_{\Delta}$. We will resolve this problem in an interesting manner in Section 6.1 when we consider the transition from FL* to a $d$-wave superconductor in the doped case.
@@
A boson with the same quantum numbers as our $B$ was introduced in earlier work [96, 68, 111], but with an important difference. In the earlier work, the expectation value of $B^{\dagger}B$ was set to be equal to the doping density. That is not the case in our work, as the doping density also includes the density of fermionic holes (see Section 2). In this section, we are considering the particle-hole symmetric case at half-filling, and $\langle B^{\dagger}B\rangle$ is non-zero even though there is no doping (this is similar to the situation in the superfluid-insulator transition of bosons [222]).
@@
These considerations lead to the energy functional $\mathcal{E}_{2}[B,U]+\mathcal{E}_{4}[B,U]$ with terms quadratic and quartic in $B$ respectively:
@@ 5 SU(2) gauge theory of the cuprate pseudogap
specified in Eqs. (2.4), (3.46), (5.3), (4.17). This Hamiltonian has 3 fermions $c_{\alpha}$, $f_{1\alpha}$, $f_{\alpha}$ whose transformations under SU(2) gauge, spin rotation, and electromagnetic charge symmetries are in Eqs. (4.8), (4.9), (4.10), with $f_{1\alpha}$ transforming just like $c_{\alpha}$. As we noted earlier, the boson $\Phi$ will be treated as a $c$-number constant for now, although we will consider its quantum and thermal fluctuations in Section 8 when we consider the transition to the FL phase at large doping. The completion of $\mathcal{H}_{\rm pseudogap}$ for the case of dynamic $\Phi$ is presented in Appendix B.
@@ 5.1 Photoemission spectrum
In the mean-field analysis of Section 2.2.4, we obtained hole pocket description of the Fermi surface in Fig. 11, along with the dispersion antinodal pseudogap in Fig. 12. Here, we will consider modification of the Fermi surface from the coupling to the bottom layers of the ALM, where the spinons reside. Should the FL* state survive at $T=0$, we expect the coupling of the hole pockets to the spinons to be relatively innocuous, as their interaction is mediated by the gapped $B$ bosons, and their zero energy modes lie at distinct momenta (see Fig. 22)—there remains a non-zero quasiparticle residue around the hole pocket at $T=0$ [77].
However, there is a stronger effect at non-zero $T$, above low $T$ phases (such as those in Fig. 23) where the spinons are confined. The computations of Pandey et al [108] of thermal fluctuations included the coupling to the spinons in $\mathcal{H}_{\rm pseudogap}$ in Eq. (5.4). This can be expected to have a stronger effect along the Brillouin zone diagonals proximate to the spinon nodes, as shown in Fig. 22. Indeed they found that upon including thermal fluctuations the backside of the hole pockets are suppressed, and spectral weight turned into Fermi arcs, similar to observations.
@@
Figure 23: From Ref. [108]. Mean-Field phase diagram of Eq. (4.17) a function of the further neighbor interactions $V_{11}$ and $V_{22}$, extending that in Fig. 18. The other parameters are r=0.732r=-0.732, $w=0.40$, $u=0$, $V_{1}=0$, $g=0.021446$, $J_{1}=K_{1}={2}/({4(1+\sqrt{2})^{2}})$. ddSC is $d$-wave superconductivity, CDWn×m is a charge density wave with a supercell with $n\times m$ lattice sites. The red cross marks the parameter values chosen for the Monte Carlo simulations of Eq. (5.5) in Ref. [108].
@@
Figure 24: The right panel shows the electron spectral function measured in photoemission computed in Ref. [108] from $\mathcal{H}_{\rm KLmf}+\mathcal{H}_{SLf}+\mathcal{H}_{Y}$ after averaging over the Monte Carlo simulation of the classical, thermal fluctuations of $B$ and $U_{{\bm{i}}{\bm{j}}}$ in Eq. (5.5). The spectrum with $B=U=0$ is in Fig. 22, showing both the hole pockets and the spinons. The left panel is experimental photoemission data from Ref. [230].
@@ 5.2 Magnetotransport
Given the role of thermal $B$ fluctuations in removing the pocket backsides in the photoemission spectrum in Fig. 24, it is legitimate to ask if the thermal $B$ fluctuations will also wash out quantum oscillations. The latter probe does not eject electrons form the sample, unlike photoemission, and so its behavior can be distinct. This question was also investigated by Pandey et al. [108]: they averaged the quantum oscillations associated with $\mathcal{H}_{\rm KLmf}+\mathcal{H}_{SLf}+\mathcal{H}_{Y}$ over gaussian fluctations of $B$. They found that the quantum oscillations are damped by thermal fluctuations but survive in regimes where the Fermi arc spectrum in Fig. 24 was well-established, as shown in Fig. 25.
@@ 6 From the pseudogap to $d$-wave superconductivity and competing orders
This section addresses the fate of the FL* pseudogap as the temperature is lowered, upon including the coupling to the ${\bm{S}}_{2}$ spin liquid in $\mathcal{H}_{\rm pseudogap}$ in Eq. (5.4). This Hamiltonian specifically choses the $\pi$-flux spin liquid of Section 3.2 for the ${\bm{S}}_{2}$ layer. We will show that for this spin liquid there is a transition to a conventional BCS-type $d$-wave superconductor, with anisotropic nodal velocities for the Boboliubov quasiparticles, and $h/(2e)$ vortices. Nevertheless the transition itself is not of the BCS type with a Cooper-pairing instability of a Fermi surface. Instead, the transition is driven by the confinement of the fractionalized excitations of the ${\bm{S}}_{2}$ spin liquid. We also find nearby instabilities to charge ordering, consistent with observations.
@@ 6.1 Anisotropic velocities in the $d$-wave superconductor
We now show how the problem of isotropic quasiparticle velocities, noted in Sections 1 and 4, is resolved by the presence of the pocket Fermi surfaces described by $\mathcal{H}_{\rm KLmf}$ in Eq. (2.4). The discussion below is based on the detailed computations presented in Refs. [113, 114].
@@ 7 The SYK model
We now turn to a different critical quantum liquid, associated with mobile electrons, the SYK model. This is a zero-dimensional model for which the absence of quasiparticle excitations is well-established. A direct and extensive application of the zero-dimensional SYK model has been to the low energy quantum theory of charged black holes, and this is reviewed elsewhere [237]. There are also connections to experiments on graphene flakes [238, 239]. But our primary interest here will be extensions to two-dimensional models relevant to the strange metal state of the cuprates and other correlated electron materials—we will turn to this in Section 8, where the zero-dimensional SYK models will motivate a theory of realistic FL*-FL transitions in two-dimensional metals. We also note other reviews [119, 120, 121] for a more in-depth discussion of the topics in Sections 7-9.
@@ 8 From the SYK model to strange metals
    where $n$ is the known conduction electron density, and $m^{\ast}$ is an electronic effective mass. In some experiments, the transport mass $m_{\rm trans}^{\ast}$ of Eq. (8.2) is used in Eq. (8.4), while other experiments use the $m^{\ast}$ determined from thermodynamic measurements. In the form Eq. (8.4), the absolute value of $\tau_{\rm trans}$ can be deduced from experimental observations. In the strange metal, such a value is found to obey ‘Planckian’ behavior with [272, 273, 274, 275, 276, 277]
    $\frac{1}{\tau_{\rm trans}}=\widetilde{\alpha}\,\frac{k_{B}T}{\hbar}\,$, (8.5)
with $\widetilde{\alpha}$ a numerical constant of order unity. Measurements of $1/\tau_{\rm trans}$ in La1.6-xNd0.4SrxCuO4 in angle-dependent magnetotransport show $\widetilde{\alpha}=1.2\pm 0.4$ [274] upon using the thermodynamic $m^{\ast}$.